OpenStudy (anonymous):

what is the antiderrivative of x^1/2

OpenStudy (anonymous):

\[x ^{1/2}\]

OpenStudy (anonymous):

2/3(x)^3/2

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

wait but i though of something different

OpenStudy (anonymous):

watch

OpenStudy (anonymous):

\[\frac{ 2x ^{3/2} }{ 3 }\]

OpenStudy (anonymous):

i thought fractions u had to flip when you brought the exponent down???

OpenStudy (anonymous):

watch

OpenStudy (anonymous):

\[x^{1/2}\]

OpenStudy (anonymous):

1/2+2/2=3/2 for the exponent

OpenStudy (anonymous):

I did flip because we are taking the antiderivative if we were taking the derivative we would then multiple that by the exponent and minus 1 from the exponent

OpenStudy (anonymous):

flim the 3/2 when you bring i down

OpenStudy (anonymous):

and then remmember th 1/ n so its 1/1/2/3=3/2

OpenStudy (anonymous):

which is wrong

OpenStudy (anonymous):



OpenStudy (anonymous):

If you divide a number by a fraction the it would be the same as multiplying that number by the reciprocal of the fraction

OpenStudy (anonymous):

ok so we just bring down the 3/2 as is in antiderrivatives

OpenStudy (anonymous):

In antiderivative we added 1 to 1/2 and made it 3/2 and then divided our x^3/2 by that 3/2 but since its a number x divided by a fraction you see its the same as multiplying by the reciprocal which is 2/3 which gives us the 2/3 x X^3/2

OpenStudy (anonymous):